OpenStudy (reneeshumpert):

The hypotenuse of a right triangle is 15 cm, and the shorter leg is 9 cm. Find the length of the other leg. 2cm cm 12 cm

OpenStudy (kidrah69):

\[a^2+b^2=c^2\]

OpenStudy (anonymous):

Do you know how to do the pythagorean theorem?

OpenStudy (kidrah69):

can you plug the numbers in?

OpenStudy (reneeshumpert):

I kinda sorta know how to do the Pythagorean theorem but it would be nice if you could put it simpler terms

OpenStudy (kidrah69):

see that those are how the sides are labled

OpenStudy (kidrah69):

The hypotenuse is always C

OpenStudy (reneeshumpert):

ok , more

OpenStudy (anonymous):

When you have a right triangle, your hypothesis will always be your c in the theorum. The legs or shorter sides will be your a and your b.

OpenStudy (kidrah69):

@Krissy3039 can finish it up dont want to confuse u :)

OpenStudy (anonymous):

So all you have to do is fill in the values you have to the equation in your case 15^2 = 9^2+b^2

OpenStudy (reneeshumpert):

ok and then what?

OpenStudy (anonymous):

You would then solve for b

OpenStudy (reneeshumpert):

so 225= 18+b^2

OpenStudy (anonymous):

almost; your value for 9^2 is wrong